# GP-MoLFormer: A Foundation Model For Molecular Generation

### v1

@@ Abstract
Transformer-based models trained on large and general purpose datasets consisting of molecular strings have recently emerged as a powerful tool for successfully modeling various structure-property relations. Inspired by this success, we extend the paradigm of training chemical language transformers on large-scale chemical datasets to generative tasks in this work. Specifically, we propose GP-MoLFormer, an autoregressive molecular string generator that is trained on more than 1.1B chemical SMILES. GP-MoLFormer uses a 46.8M parameter transformer decoder model with linear attention and rotary positional encodings as the base architecture. We explore the utility of GP-MoLFormer in generating novel, valid, and unique SMILES. Impressively, we find GP-MoLFormer is able to generate a significant fraction of novel, valid, and unique SMILES even when the number of generated molecules is in the 10 billion range and the reference set is over a billion. We also find strong memorization of training data in GP-MoLFormer generations, which has so far remained unexplored for chemical language models. Our analyses reveal that training data memorization and novelty in generations are impacted by the quality of the training data; duplication bias in training data can enhance memorization at the cost of lowering novelty. We evaluate GP-MoLFormer’s utility and compare it with that of existing baselines on three different tasks: de novo generation, scaffold-constrained molecular decoration, and unconstrained property-guided optimization. While the first two are handled with no additional training, we propose a parameter-efficient fine-tuning method for the last task, which uses property-ordered molecular pairs as input. We call this new approach pair-tuning. Our results show GP-MoLFormer performs better or comparable with baselines across all three tasks, demonstrating its general utility for a variety of molecular generation tasks.
@@
Identifying molecules with desirable properties from the vast landscape of possibilities is daunting. It has been postulated that only $10^{8}$ out of $10^{23}$$10^{60}$ possible small drug-like molecules are synthesizable [1]. As the search space is enormous and high-throughput screening of molecules is costly and time-consuming, this requires a thorough understanding of the chemical data manifold. Recently, similar to what has been experienced in computer vision and natural language processing, deep generative models have made great strides in modeling molecular distributions and sampling new molecules from them in de novo or targeted manners. Among those efforts, a significant fraction use string-based representations of molecules as the input; thus, techniques explored in language modeling, such as causal and masked language modeling, are becoming widely used in building molecular deep neural models.
@@
Interestingly, much of the recent performance gains for natural language models have come from scale — in terms of the number of parameters and the number of training samples [2, 3, 4]. It is also clear that larger language models that can memorize training data show improved generalization [5, 6, 7]. However, perhaps surprisingly, relatively less work has been done in understanding the impact of scale and the impact of training data memorization on the performance of generative models of molecules. Specifically, it remains under-explored to what extent a causal large language model of molecules, trained on large-scale (>100M) training data, memorizes its training data and demonstrates such memorization in its generations.
@@
In the context of memorization of language models, it is now known that data that is seen during training many times is memorized more, and de-duplication of training data plays a big role in preventing such memorization  [5, 6, 8]. In chemical language modeling tasks, molecules in training data originate from publicly available databases such as ZINC [9] and PubChem [1]. It is known that certain molecules, as well as certain molecular features, are over-represented in those databases [10]. How such bias is perpetuated by generative chemical language models remains relatively unknown.
@@
To bridge these gaps, in this work, we present a family of generative pre-trained molecular foundation models for the unconstrained and targeted generation of novel molecules, which are based on the recently published Molecular Language transFormer (MoLFormer) architecture [11]. We refer to these Generative Pre-trained models as GP-MoLFormer. The base transformer architecture of our decoder-only GP-MoLFormer, consisting of $\approx 47$M parameters, uses an efficient linear attention mechanism, together with rotary positional encodings, analogously to MoLFormer [11]. The model is then trained with a causal language modeling objective on a large corpus of 0.65–1.1 billion canonicalized SMILES strings of small molecules from publicly available chemical databases. We evaluate memorization in GP-MoLFormer, in the light of the bias present in training data, and how such memorization affects the model’s generation performance in the de novo molecule generation task. To our knowledge, this is the first report on effect of training data memorization in a generative pre-trained chemcial language model. Experiments demonstrate that GP-MoLFormer is able to generate novel, unique, diverse, and valid molecules even when the generation pool reaches a size of 10B while showing consistent memorization of training data. We further explore how representational bias encoded in the public chemical databases is perpetuated by a generative chemical language model and is reflected in its generations.
@@
We further extend GP-MoLFormer to two targeted molecular design tasks: scaffold-constrained molecular decoration and unconstrained property-guided optimization. For the first, we exploit GP-MoLFormer’s causal language modeling ability and establish GP-MoLFormer’s ability to handle the task without undergoing any task-specific tuning. For the second, we provide a prompt-tuning or soft prompt-learning algorithm that learns from partial orderings of molecules. We name this method pair-tuning. Results show that pair-tuning on GP-MoLFormer provides on par or better performance in three different property optimization tasks, namely (i) drug-likeness optimization, (ii) penalized logP optimization, and (iii) optimization of dopamine type 2 receptor binding activity.
@@
GP-MoLFormer Framework
We construct the GP-MoLFormer architecture (Figure 1A) using a 12-layer linear attention transformer with 12 attention heads and a hidden state size of 768 and generalized random Fourier feature maps [12] for the linear attention. The input representation considered is chemical SMILES. We choose a relative positional embedding, namely rotary positional embedding [13], to model the token dependency within SMILES better using rotary attention. During training, GP-MoLFormer uses a causal modeling objective of predicting the next token given the context history of prior tokens in the input SMILES strings. Given the size of the model and the efficient linear attention, GP-MoLFormer takes only around 3 milliseconds for a single forward pass during generation. After pre-training, we quantitatively assess the performance of GP-MoLFormer on the generation of de novo molecules before applying a novel prompt-tuning algorithm for molecular optimization (Figure 1B).
@@
The task under consideration is to generate random molecules by sampling from the generative model. As downstream optimization can rely on these randomly generated molecules as starting points, the generated distribution must contain novel, diverse, and unique molecules.
Effect of training dataset and generation pool
The generative ability of a chemical language model can be affected by the scale of the training data. Further, bias in training data can also contribute to the quality of generation. To disentangle these factors, we report GP-MoLFormer’s performance on two independently varying dimensions: First, we vary the size and quality of pre-training data. For this dimension, we consider training the GP-MoLFormer on (i) 1.1B SMILES extracted from ZINC and PubChem, as described in Ross et al. (2022)[11], and (ii) a de-duplicated 650M subset of (i). GP-MoLFormer trained on (i) is referred hereafter GP-MoLFormer-Raw, while GP-MoLFormer trained on (ii) is termed as GP-MoLFormer-Clean. Secondly, we analyze the effect of varying the number of generated molecules from 30k to 10B molecules. It should be noted that existing molecular generation benchmarks consider reporting performance metrics for a relatively small 10–30k generations, as those models are designed to target a specific distribution of molecules, e.g., synthetic molecules with biological activity or natural products, and are trained on 1–100M samples[14]. In contrast, GP-MoLFormer is trained on 650M–1.1B SMILES and the dataset captures the relative abundance of molecules, as well as the presence of the same molecule in different context, as found in chemical databases.
We report in Table 1 the percentage of novel (unseen in training), valid (syntactically correct), and unique (not previously generated) molecules for both GP-MoLFormer-Raw and GP-MoLFormer-Clean, for generation size of 30k to 10M. The results suggest that the fraction of novel generations stays at a consistent $\approx$ 32% for GP-MoLFormer-Raw when the number of total generated molecules is within this range. Novelty in GP-MoLFormer-Clean is $\approx$7–8higher compared to that of GP-MoLFormer-Raw for all generation pool sizes and rises with the increasing pool size. GP-MoLFormer, irrespective of training data, outputs chemically valid SMILES almost all the time. While the percentage of unique molecules drops slightly with increasing generation pool size, it still is over 99% for 10M generations. This result confirms that (i) GP-MoLFormer memorizes training data, as seen from the high number of exact matches ($1-$novelty, which can be up to 60%) with training molecules; and (ii) memorization is less when the training data is de-duplicated, enabling more novel generation. Nevertheless, in all cases studied here, GP-MoLFormer is capable of generating novel, diverse, and valid molecules.
We continue exploring the quality of generations while increasing the number of samples generated to billions for GP-MoLFormer-Raw. Results are shown in Table 2. At or beyond 1B generations, the fraction of novel and unique generations drops but still remains significant. Even for 10B generations, GP-MoLFormer-Raw is able to generate a significant 16.7% novel molecules. And all throughout, 99% of generations are chemically valid.
Training Size = 650M Training Size = 1.1B
Generation Size Novel Unique Valid Novel Unique Valid
30k 0.390 0.997 1.000 0.323 0.997 0.997
100k 0.393 0.996 0.999 0.326 0.998 0.998
1M 0.395 0.996 0.999 0.323 0.996 0.997
10M 0.400 0.991 0.994 0.322 0.989 0.997
Table 1: Novelty, validity, and uniqueness of different number of generations for GP-MoLFormer trained with 650M (Clean) and 1.1B (Raw) training set.
Size Novel Unique Valid
30k 0.323 0.997 0.997
100k 0.326 0.998 0.998
1M 0.323 0.996 0.997
10M 0.322 0.956 0.997
100M 0.327 0.989 0.997
1B 0.278 0.611 0.997
10B 0.167 0.223 0.997
Table 2: Metrics on GP-MoLFormer-Raw generations of different scale.
Effect of sampling temperature
We also report the generative performance of GP-MoLFormer-Raw at several softmax temperature values, as we use multinomial sampling during decoding to encourage novel generation. In fact, recent work has reported the utility of multinomial sampling over beam search for generating molecules that are simultaneously novel and useful (in terms of synthesis and biological testing) [15]. Results are shown in Figure 2, which shows that at around $T=1.2$, GP-MoLFormer-Raw performs the best in terms of balancing high novelty (47%), uniqueness (99%), and validity (98%). Further, when we compare novelty of the same generated sets with respect to a random subset of 130M molecules from training set, we see novelty is significantly higher (around 90% at $T=1.2$), showing the effect of training data scale on generation novelty — there are simply more opportunities to match with a larger reference set. All results elsewhere in this paper are generated using $T=1.0$.
Figure 2: Effect of temperature during multinomial sampling of 100k generated molecules on the percentage of novel molecules (blue) with respect to full 1.1B training set, novelty with respect to a random 130M subset of the training (orange), validity (green), and uniqueness (red) of generations from GP-MoLFormer-RAW.
Discussions on the effect of training data bias on generations
Within the Raw training set, a notable 45% of SMILES strings were found duplicates. Some of this is due to molecules that appear multiple times as different isomers, which are pre-processed into the same SMILES representation in our canonicalization pipeline. There is also some overlap between the two databases — ZINC and PubChem. This popularity of certain molecules may or may not be related to true factors, such as the molecule being useful in a multitude of applications because of its synthetic ease or lower cost, or a combination of all of those. It is reported that such skewed distribution can also originate from anthropogenic attention bias [10], where some molecules are studied extensively because scientists themselves or their peers “like” them. Nevertheless, such overrepresenation of certain molecules, either originating from human cognitive biases, heuristics, social influences, from rationally made choices, from multiple isomeric instances, or a combination of all of the above, can lead to bias in databases.
Data de-duplication is the first step towards removing such bias, which reduces the concentration of the high density regions of the data manifold. In this case, de-duplication removes isomeric population information as well as repeated molecules across databases. The de-duplicated data is closer to a data manifold that has more homogenized density all over. Training on such a data manifold results in higher novelty in generations, as found for GP-MoLFormer-Clean when compared to GP-MoLFormer-Raw, as shown in Table 1.
While existing molecular generative models trained on a much smaller and much focused datasets have demonstrated near-perfect (100%) novelty in generations, they are for most part is not suitable for studying the trade-off between training data memorization and generation novelty. Investigating such phenomenon requires studying a generative chemical (language) model that has been trained on a broader-purpose and much larger dataset at scale. Our experiments attempt to address this under-explored aspect in this study. As shown in Table 1, novelty in GP-MoLFormer generations are lower compared to $\approx 100\%$ reported by baselines[14], but still sufficiently high for practical use. When compared with recent baselines, GP-MoLFormer generations are more dissimilar to test molecules (Table 3), though GP-MoLFormer’s test set is more diverse. And, finally, the low novelty in GP-MoLFormer’s generations is reflective of modeling its training set that represents the relative usage of molecules in real-world.
Similarly, the present study highlights the importance of studying generated sets of different sizes to obtain a comprehensive view of the quality of generations, particularly when the generative model is trained on data at scale. As GP-MoLFormer-Clean aims to capture a training data manifold of more uniform density, which is enabled by de-duplicating the training SMILES, we see a 1% rise in novelty as we increase the number of generated samples from 30k to 10M. A similar observation has been reported in image generation [16] and language generation [17, 8]. To summarize, novelty in generations is influenced by the support provided by both the training distribution and the generated distribution, and should therefore be assessed relative to the sizes and diversity of those two sets.
These results in Table 1 complement and support earlier efforts focusing on studying scaling behaviors of chemical language models. One such noteworthy effort along this line is Frey et al. (2023)[18], where neural-scaling behaviour in large chemical models was investigated by studying models with over 1B parameters. However, the models tested in that work were only pre-trained on datasets of size up to $\approx$ 10M data points, which is very small compared to the size of the chemical universe. In Ross et al. (2022)[11], the scaling behavior of MoLFormer, which is a transformer-based molecular encoder built using a masked language modeling objective, was studied. That work clearly established the scaling behavior underlying adaptation of a pre-trained model across downstream tasks, in which the number of model parameters was up to 47M while the number of training points considered was >1B. It was shown that a MoLFormer trained on 100M SMILES consistently underperformed across a wide variety of property prediction tasks, including quantum mechanical and physiological, when compared to the model trained on >1.1B SMILES, indicating predictive ability may benefit from such bias in training data. In contrast, the results in Table 1 show that a generative chemical language model trained on cleaner de-duplicated training data produces more novel generations.
Comparison of GP-MoLFormer with baselines
@@
Next, we compared GP-MoLFormer-Clean with different baseline models, such as the character-level recurrent neural network (CharRNN) [14], SMILES variational autoencoder (VAE) [14], junction tree VAE (JT-VAE) [19], latent inceptionism on molecules (LIMO) [20], and MolGen-7b [21]. We consider baselines that are not fine-tuned to particular molecular domains. Except MolGen-7b, all baselines were trained and tested on datasets from MOSES[14], whose origin is the ZINC Clean Leads dataset[9]. The size of that training set is 1.6M. MolGen-7b was trained on 100M filtered molecules from ZINC-15[9] as detailed in Irwin et al. (2022)[22]. LIMO and MolGen-7b are trained using an alternative molecular string representation, SELFIES [23], that guarantees 100% validity of generated molecules. We, in contrast, train GP-MoLFormer on SMILES as recent work shows that training a generative language model on SELFIES may hurt model’s exploratory ability [24]. All baseline performances are reported on their corresponding test set consisting of 175k molecules (if the original test set was larger, this is a randomly selected subset). While GP-MoLFormer-Clean’s performance is estimated on a held-out test set that is of similar size, we found this test set to be more diverse in terms of number of unique scaffolds present within the set (126k compared to 124k in the ZINC-15 subset and 77k in the MOSES set) and by comparing different property distributions with that of the other baselines (see Figure 3).
@@
Standard metrics for evaluating model-generated molecules are reported in Table 4 for a generation set of 30k molecules. When compared to baselines, GP-MoLFormer-Clean is equally performant in generating unique, valid, and novel molecules that share high cosine similarity with the corresponding reference molecules at the fragment (Frag) level, consistent with low Fréchet ChemNet Distance (FCD)[25]. At the same time, GP-MoLFormer-Clean generates molecules with high internal diversity (IntDiv), i.e., average pairwise dissimilarity. The scaffold cosine similarity (Scaf) and similarity to the nearest neighbor in the test set (SNN) of GP-MoLFormer-Clean is comparable to that of baselines for 30k generations. All metrics are computed using MOSES[14].
@@
In order to quantify the amount of novelty in generations, we also calculated the pairwise similarity between novel and unique generations and molecules from the corresponding 175k sample test set. We then report both the average similarity per generated molecule and the maximum similarity per generated molecule over the test set. These results are presented in Table 3. GP-MoLFormer-Clean is slightly lower than MolGen-7b in both average mean and average maximum similarity, indicating generations that are more dissimilar with respect to its test set. LIMO is much lower than both of these, though, as we see in Table 4, the outputs of this model do not match its test set well, so this is to be expected. Taken together, this means GP-MoLFormer-Clean is able to generate molecules that are about as novel as other models, despite the seemingly low novelty scores in Table 1 (see also, the discussion of Figure 2).
@@
GP-MoLFormer-Clean 0.1354 0.4533
Table 3: Average Tanimoto similarity between 30k randomly generated molecules and all respective test molecules (mean) or most similar test molecule (max).
We next investigate how these metrics change with varying number of generated and test molecules. Table 5 shows that, with increasing the generated pool size, scaffold similarity with respect to the test molecules becomes $>0.9$ while SNN reaches $>0.5$ when compared against 175k held-out test samples. When a larger test set of 1M molecules is used, further increases in both scaffold similarity and SNN are observed. These results suggest that GP-MoLFormer is able to model the training data faithfully. Further, our work implies that, with increasing diversity of the training data, the typical metrics used in assessing molecular generative models, such as various similarity measures with respect to a test set, should be analyzed with generation and test sets that are larger in size compared to what is typically used in the field.
Figure 3: Property distributions of different test datasets — MOSES, ZINC-15 (MolGen-7b), and GP-MoLFormer-Clean (ours). Clockwise from top left: octanol-water partition coefficient, drug-likeness, synthetic accessibility, molecular weight. Our test distributions are consistently wider (more diverse) than the other baselines.
[…]
CharRNN [14] 0.9998 0.9242 0.6015 0.8562 0.0732
VAE [14] 0.9984 0.9386 0.6257 0.8558 0.0990
JT-VAE [19] 0.9965 0.8964 0.5477 0.8551 0.3954
LIMO [20] 0.6989 0.0079 0.2464 0.9039 26.78
MolGen-7b [21] 0.9999 0.6538 0.5138 0.8617 0.0435
GP-MoLFormer-Clean 0.9998 0.7383 0.5045 0.8655 0.0591
Table 4: Comparison of 30k generations with a held-out test set of size 175k. Baseline performances for CharRNN, VAE, and JT-VAE are taken from MOSES[14]. LIMO is reproduced using their random generation model trained on the MOSES data. These models all use the default MOSES test split for reference. The MolGen-7b baseline uses the pre-trained model from Hugging Facehttps://huggingface.co/zjunlp/MolGen-7b with multinomial sampling ($T=1.0$) and is tested on a random 175k subset of the original test data. GP-MoLFormer-Clean is tested with respect to a held-out 175k set from its training data.
Test size Gen. size Frag Scaf SNN IntDiv FCD
175k 30k 0.9998 0.7383 0.5045 0.8655 0.0591
100k 0.9998 0.8653 0.5045 0.8657 0.0279
1M 0.9998 0.9375 0.5040 0.8658 0.0178
1M 30k 0.9998 0.7702 0.5738 0.8655 0.0646
100k 0.9998 0.9026 0.5740 0.8657 0.0331
1M 0.9998 0.9786 0.5739 0.8658 0.0227
Table 5: Investigation on the interplay between generation set size and test set size on modeling the molecular distribution.
@@
We further subject GP-MoLFormer to the task of scaffold-constrained generation. For this experiment, we first take the five unique scaffolds from the dopamine receptor $\mathrm{D}_{2}$ (DRD2) active binder dataset validation split[26]. These scaffolds contain between two to four attachment points. We perform a pre-processing step for each unique scaffold, generating every possible randomized SMILES representation of that scaffold. Then we sort the resulting candidates according to the distance of the “*” characters, also known as the attachment point, to the end of the string. For multiple *s, we sum the distances. As an example, C1(=O)N(CCN1*)* would score $2(2+0)$ while C1N(*)C(=O)N(*)C1 would score $15(12+3)$. If multiple representations are equivalently optimal, we save all of them. During the generation step, we provide the candidates produced in this pre-processing step as input to GP-MoLFormer.

### v2

@@ Abstract
Transformer-based models trained on large and general purpose datasets consisting of molecular strings have recently emerged as a powerful tool for successfully modeling various structure-property relations. Inspired by this success, we extend the paradigm of training chemical language transformers on large-scale chemical datasets to generative tasks in this work. Specifically, we propose GP-MoLFormer, an autoregressive molecular string generator that is trained on more than 1.1B (billion) chemical SMILES. GP-MoLFormer uses a 46.8M parameter transformer decoder model with linear attention and rotary positional encodings as the base architecture. GP-MoLFormer’s utility is evaluated and compared with that of existing baselines on three different tasks: de novo generation, scaffold-constrained molecular decoration, and unconstrained property-guided optimization. While the first two are handled with no additional training, we propose a parameter-efficient fine-tuning method for the last task, which uses property-ordered molecular pairs as input. We call this new approach pair-tuning. Our results show GP-MoLFormer performs better or comparable with baselines across all three tasks, demonstrating its general utility for a variety of molecular generation tasks. We further report strong memorization of training data in GP-MoLFormer generations, which has so far remained unexplored for chemical language models. Our analyses reveal that training data memorization and novelty in generations are impacted by the quality and scale of the training data; duplication bias in training data can enhance memorization at the cost of lowering novelty. We further establish a scaling law relating inference compute and novelty in generations.
@@
Identifying molecules with desirable properties from the vast landscape of possibilities is daunting. As the search space is enormous and high-throughput screening of molecules is costly and time-consuming, this requires a thorough understanding of the chemical data manifold. Recently, similar to what has been experienced in computer vision and natural language processing, deep generative models have made great strides in modeling molecular distributions and sampling new molecules from them in de novo or targeted manners. Among those efforts, a significant fraction use string-based representations of molecules as the input; thus, techniques explored in language modeling, such as causal and masked language modeling, are becoming widely used in building molecular deep neural models.
@@
Interestingly, much of the recent performance gains for natural language models have come from training at scale — in terms of the number of parameters and the number of training samples [1, 2, 3]. It is reported that larger language models that can memorize training data show improved generalization [4, 5, 6]. Furthermore, data that is seen during training many times is memorized more and de-duplication of training data plays a big role in preventing such memorization  [4, 5, 7, 8].
@@
However, less work has been done in understanding the impact of training data scale and its memorization on the performance of generative models of molecules. Specifically, it remains under-explored to what extent a causal large language model of molecules, trained on large-scale (>100M) training data, memorizes its training data and demonstrates such memorization in its generations. In chemical language modeling tasks, molecules in training data originate from publicly available databases such as ZINC [9] and PubChem [10]. It is known that certain molecules as well as certain molecular features are over-represented in those databases [11] but how such training bias is perpetuated by generative chemical language models remains relatively unknown.
An additional dimension of scaling in traditional large language models that has been investigated recently is inference-time compute scaling [12]. It has been shown that with increasing inference compute, performance across multiple tasks can be increased for the same model, as it allows better coverage of the search space. On the other hand, the effect of inference scaling by increasing number of generations is under-explored for molecular generative models.
@@
To bridge these gaps, in this work, we present a family of generative pre-trained molecular foundation models for the unconstrained and targeted generation of novel molecules. These decoder-only models are based on the recently published Molecular Language transFormer (MoLFormer) architecture [13]. We refer to these Generative Pre-trained models as GP-MoLFormer. The base transformer architecture of our GP-MoLFormer consists of $\approx 47$M parameters and uses an efficient linear attention mechanism together with rotary positional encodings — analogous to MoLFormer [13] but using decoder instead of encoder blocks (Figure 1A). The model is then trained with a causal language modeling objective on a large corpus of 0.65–1.1 billion canonicalized SMILES strings of small molecules from publicly available chemical databases.
@@
We evaluate GP-MoLFormer on an unconditional de novo generation task as well as to two targeted molecular design tasks: scaffold-constrained molecular decoration and unconstrained property-guided optimization. For scaffold decoration, we exploit GP-MoLFormer’s causal language modeling ability and establish GP-MoLFormer’s ability to handle the task without undergoing any task-specific tuning. For the optimization task, we provide a prompt-tuning or soft prompt-learning algorithm that learns from partial orderings of molecules. We name this method pair-tuning (Figure 1B). Results show that pair-tuning on GP-MoLFormer provides on par or better performance in three different property optimization tasks, namely (i) drug-likeness optimization, (ii) penalized logP optimization, and (iii) optimization of dopamine type 2 receptor binding activity.
We further extensively evaluate quality of GP-MoLFormer-generated molecules, in the light of the training data scale and the bias present in the training data. Experiments reveal significant memorization in de novo generations affecting novelty therein. We further analyze how representational bias encoded in the public chemical databases is perpetuated by a generative chemical language model and is reflected in its generation quality. To our knowledge, this is the first report on effect of training data memorization in a generative pre-trained chemical language model. Further, we investigate the effect of inference compute as another scaling dimension by increasing the number of generated samples and establish a inference scaling law relating number of generations with novelty in them. Experiments demonstrate that novelty in de novo generations by GP-MoLFormer drops when number of generated samples reaches a scale of $\approx$ 1B. Nevertheless, GP-MoLFormer is able to generate novel, unique, diverse, and valid molecules even when the generation pool reaches a size of 10B, while showing consistent memorization of training data.
Our main contributions are:
  • We present an autoregressive transformer-based SMILES decoder, GP-MoLFormer, trained on up to 1.1 billion chemical SMILES.
  • We report the performance of GP-MoLFormer models on de novo and scaffold-constrained generation tasks.
  • We provide a parameter-efficient finetuning method, which utilizes property-ranked molecule pairs as input, for property-guided molecule generation and show its effectiveness on three different tasks.
  • We further study training and inference scaling of GP-MoLFormer and report the effect on generation novelty. We reveal that training data duplication significantly affects novelty in generations. We also report a scaling behavior relating inference compute and novelty that follows exponential decay, while showing that GP-MoLFormer can generate a notable fraction of novel SMILES, even when number of generations reaches 10B.
  • We demonstrate higher novelty and diversity in GP-MoLFormer generations compared to baselines, which is attributed to its training data scale.
@@
GP-MoLFormer uses a causal modeling objective of predicting the next token given the context history of prior tokens in the input SMILES strings. For details of model architecture and training, see the Methods section. After pre-training, we quantitatively assess the performance of GP-MoLFormer on de novo molecule generation and scaffold-constrained molecule decoration tasks, before applying a novel prompt-tuning algorithm for molecular optimization.
@@
The task under consideration is to generate random, valid molecules by sampling from the generative model. As downstream optimization may rely on these randomly generated molecules as starting points, the generated distribution must contain novel, diverse, and unique molecules. Here we compare a GP-MoLFormer model trained on 650M unique SMILES, that is a de-duplicated subset of the 1.1B SMILES extracted from ZINC and PubChem, as described in Ross et al. (2022)[13]. This model variant is referred to as GP-MoLFormer-Uniq hereafter.
@@
We compare GP-MoLFormer-Uniq with different baseline models, such as the character-level recurrent neural network (CharRNN) [14], SMILES variational autoencoder (VAE) [14], junction tree VAE (JT-VAE) [15], latent inceptionism on molecules (LIMO) [16], and MolGen-7b [17]. Except MolGen-7b, all baselines were trained and tested on datasets from MOSES[14], whose origin is the ZINC Clean Leads dataset[9]. The size of that training set is 1.6M. MolGen-7b was trained on 100M filtered molecules from ZINC-15[9] as detailed in Irwin et al. (2022)[18]. LIMO and MolGen-7b are trained using an alternative molecular string representation, SELFIES [19], that guarantees 100% validity of generated molecules. We, in contrast, train GP-MoLFormer-Uniq on SMILES as recent work shows that training a generative language model on SELFIES may hurt model’s exploratory ability [20]. All baseline performances are reported on their corresponding test set consisting of 175k molecules (if the original test set was larger, this is a randomly selected subset).
First, we note that GP-MoLFormer-Uniq (and GP-MoLFormer) exhibits excellent validity and uniqueness at standard generation size (30/10k). See Supplementary Table S1 for comparison with the baseline models. At the same time, we argue that these metrics are insufficient to measure generation at scale. Furthermore, as we show later, novelty is dependent on training set size in addition to generation size so models trained on different size datasets are not directly comparable. See the Scaling results section below for further discussion.
@@
Standard metrics for evaluating model-generated molecules are reported in Table 1 for a generation set of 30k molecules. When compared to baselines, GP-MoLFormer-Uniq is equally performant in generating molecules that share high cosine similarity with the corresponding reference molecules at the fragment (Frag) level, consistent with low Fréchet ChemNet Distance (FCD)[21]. At the same time, GP-MoLFormer-Uniq generates molecules with high internal diversity (IntDiv), i.e., average pairwise dissimilarity. The scaffold cosine similarity (Scaf) and similarity to the nearest neighbor in the test set (SNN) of GP-MoLFormer-Uniq is comparable to that of baselines for 30k generations. All these metrics are computed using the MOSES[14] framework (we limit our scope to MOSES in this study, although we note that myriad other benchmarks are available for evaluating generative molecular models[22, 23, 24]).
We further report analogous metrics computed using MoLFormer [13] embeddings as the chemical features and estimate distances using those embeddings as a measure of similarity (under column MoLFormer-based metrics; see Table 1 caption for details). The trends observed on these metrics further support the fact that GP-MoLFormer-Uniq generates a molecular distribution that is close to the training in terms of fragment and scaffold composition as well as projections to MoLFormer space, while exhibiting high diversity, when compared to baselines.
MOSES metrics MoLFormer-based metrics
[…]
CharRNN [14] 0.9998 0.9242 0.6015 0.8562 0.0732 5.735 13.03 0.1515
VAE [14] 0.9984 0.9386 0.6257 0.8558 0.0990 5.549 13.09 0.2531
JT-VAE [15] 0.9965 0.8964 0.5477 0.8551 0.3954 6.312 12.97 1.700
LIMO [16] 0.6989 0.0079 0.2464 0.9039 26.78 11.41 13.08 162.0
MolGen-7b [17] 0.9999 0.6538 0.5138 0.8617 0.0435 6.788 12.58 0.1237
GP-MoLFormer-Uniq 0.9998 0.7383 0.5045 0.8655 0.0591 6.970 13.10 0.1844
Table 1: Comparison of 30k generations with a held-out test set of size 175k. “MOSES metrics” columns refer to the typical set of generation performance metrics (see Polykovskiy et al.[14] for details) computed with respect to the following sets: Baseline performances for CharRNN, VAE, and JT-VAE are taken from MOSES[14]. LIMO is reproduced using their random generation model trained on the MOSES data. These models all use the default MOSES test split for reference. The MolGen-7b baseline uses the pre-trained model from Hugging Facehttps://huggingface.co/zjunlp/MolGen-7b with multinomial sampling ($T=1.0$) and is tested on a random 175k subset of the original test data. GP-MoLFormer-Uniq is tested with respect to a held-out 175k set from its training data. “MoLFormer-based metrics” columns refer to analogous metrics computed using MoLFormer [13] embedding distances instead of Tanimoto similarity. DNN is the average Euclidean distance from generated molecules to the nearest molecule from the test set. IntDiv2 is the average pairwise Euclidean distance between generated molecules. FMD is the Fréchet distance between the MoLFormer embedding distributions. Bold values indicate the best model for a given metric.
@@
We also calculated the pairwise Tanimoto similarity between novel and unique generations and molecules from the corresponding 175k sample test set using molecular fingerprints as features. We then report both the average similarity per generated molecule and the maximum similarity per generated molecule over the test set. These results are presented in Table 2. GP-MoLFormer-Uniq is slightly lower than MolGen-7b in both average mean and average maximum similarity, indicating generations are slightly more dissimilar with respect to its test set. LIMO results are much lower than both of these, though, as we see in Table 1, the outputs of this model do not match its test set well, so this is to be expected. We also provide discussion of scaffold novelty in Supplementary Figure S1. Taken together, these analyses mean that GP-MoLFormer-Uniq is able to generate molecules that are distinct at the scaffold level as well as at the fingerprint level, to a similar extent to other models, despite the seemingly low string level novelty scores in the Supplementary Table S1 (see also the discussion of the Supplementary Figure S2).
@@
GP-MoLFormer-Uniq 0.1354 0.4533
Table 2: Average Tanimoto similarity between 30k randomly generated molecules and all respective test molecules (mean) or most similar test molecule (max). Bold values indicate the highest similarity to test molecules.
Figure 2 shows the property distributions of the different test sets, as well as of molecules generated using GP-MoLFormer-Uniq. The generated distribution shows very good reproduction of the corresponding test distribution. Furthermore, while GP-MoLFormer-Uniq’s performance is estimated on a held-out test set that is of similar size, we found this test set to be more diverse in terms of number of unique scaffolds present within the set (126k compared to 124k in the ZINC-15 subset and 77k in the MOSES set) and by comparing different property distributions with that of the other baselines. More analyses on how these statistics change with training data variations and generated pool size can be found later (see Discussions).
We also examine the domain adaptation of GP-MoLFormer via down-stream finetuning on a set of 36.7M drug-like molecules from PubChem [10]. In Figure 2, we show results of this fine-tuned model, referred as GP-MoLFormer-Druglike. The fine-tuning set contains molecules with QED $>0.6$ [25]. Results show that, the generated molecules undergo a distribution shift in properties as expected. For example, the QED distribution is shifted toward right compared to GP-MoLFormer-Uniq. We also provide examples of de novo generated molecules in the Supplementary Figure S3.
Figure 2: Property distributions of different test datasets — MOSES, ZINC-15 (MolGen-7b), and GP-MoLFormer-Uniq (ours) — along with generated samples from GP-MoLFormer-Uniq and GP-MoLFormer-Druglike. Clockwise from top left: octanol-water partition coefficient, drug-likeness, synthetic accessibility, molecular weight. Our test distributions are consistently wider (more diverse) than the other baselines. Furthermore, the generated distribution matches the corresponding test distribution almost exactly. In comparison to GP-MoLFormer-Uniq, a density shift toward higher QED values with GP-MoLFormer-Druglike can be observed, as expected.
@@
We further subject GP-MoLFormer to the task of scaffold-constrained generation. For this experiment, we utilize the GP-MoLFormer trained on 1.1B SMILES. We first take the five unique scaffolds from the dopamine receptor $\mathrm{D}_{2}$ (DRD2) active binder dataset validation split[26]. These scaffolds contain between two to four attachment points. We perform a pre-processing step for each unique scaffold, generating every possible randomized SMILES representation of that scaffold. Then we sort the resulting candidates according to the distance of the “*” characters, also known as the attachment point, to the end of the string. For multiple *s, we sum the distances. As an example, C1(=O)N(CCN1*)* would score $2(2+0)$ while C1N(*)C(=O)N(*)C1 would score $15(12+3)$. If multiple representations are equivalently optimal, we save all of them. During the generation step, we provide the candidates produced in this pre-processing step as input to GP-MoLFormer.